OpenStudy (anonymous):

What is the slope of the line whose equation is 6x +3y = 9?

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

@spoprockstar

OpenStudy (anonymous):

https://www.desmos.com/calculator/stww6q3gta

OpenStudy (anonymous):

6x + 3y = 9 We have to rearrange to slope-intercept form. FIrst we subtract 6x to both sides: 3y = -6x + 9 Now we divide 3 to both sides: y = -2x + 3 Now it's in the form of y = mx + b, where 'm' is the slope, so what's the slope here?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

No, that's wrong. y = -2x + 3 y = mx + b 'm' is slope, so slope here is?

OpenStudy (anonymous):